OpenStudy (anonymous):

solve 3x+2/x-1=3x+4/x+1

OpenStudy (anonymous):

x = -3/2

OpenStudy (anonymous):

Common Denomiator:(x+1)(x-1)

OpenStudy (anonymous):

((x+1)(3x+2))/(x-1)(x+1) = (3x+4)(x-1)/(x+1)(x-1) (x+1)(3x+2) = (3x+4)(x-1) Now, solve.

OpenStudy (anonymous):

x = -1

OpenStudy (anonymous):

x = -3/2

OpenStudy (anonymous):

my fault i did it backwards... x = -1.5 or -3/2

OpenStudy (anonymous):

it is so siple dude..just cross multiply u will got a quadratic equation..and just solve that